Evaluate: 11C6
1. 66
2. 120
3. 24
4. 462
Is there someone that can help me with this, please? Thank you.

Answer by MathLover1(20855)


11C6=11!/(11-6)!*6! = (11*10*9*8*7*6)/(1*2*3*4*5*6) = 462